OpenStudy (anonymous):

Assume that the box contains 11 balls: 4 red, 4 blue, and 3 yellow. As in the text, you draw one ball, note its color, and if it is yellow replace it. If it is not yellow you do not replace it. You then draw a second ball and note its color. (1) What is the probability that the second ball drawn is yellow? (2) What is the probability that the second ball drawn is red?

OpenStudy (anonymous):

(1) if the first ball was yellow then the answer is 2/9 if the first ball wasnt yellow then the answer is 2/8 (2)if the first ball was yellow then the answer is 5/9 if the first ball wasnt yellow then the answer is 1/8 OR 2/8 (because the first ball can be red OR blue) ;) hope this helps!

OpenStudy (cwrw238):

if it is replaced then P(2nd = yellow) = 3/11 not replaced .. .. .. = 3/10

OpenStudy (anonymous):

im confused what is the answer for part one

OpenStudy (anonymous):

2/9

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

its not?

OpenStudy (cwrw238):

i'm confused also how did u get 2/9 kelly?

OpenStudy (anonymous):

nope my homework thing tells you if its wrong

OpenStudy (anonymous):

homework thing? is it multiple choice?

OpenStudy (cwrw238):

i think its 3/10 + 3/11 = 63/110

OpenStudy (anonymous):

no it just tells you if its right or wrong

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

63/110 isn't right

OpenStudy (anonymous):

thats not wat i wrote..?

OpenStudy (cwrw238):

hhmm - must admit probability not my strong point

OpenStudy (anonymous):

yah.. i guess its not really mine either..;)

OpenStudy (anonymous):

ok, well, it looked right... we can try to pick it apart if you're sure it's not right.

OpenStudy (cwrw238):

oh - didn't include probabilities oif first draw

OpenStudy (anonymous):

Prob of yellow on 2nd ball: P of yellow on first ball = 3/11, then replace, then 3/11 on 2nd ball and P of not yellow on first ball = 8/11, then no replace, then yellow 3/10 on 2nd ball

OpenStudy (anonymous):

(properly, that middle "and" should be an "or")

OpenStudy (anonymous):

(3/11)(3/11) + (8/11)(3/10) = 9 /121 + 24/110 Maybe? I'm thinking and typing... this isn't my best topic either...

OpenStudy (cwrw238):

yellow first and second = 3/11 * 3/11 yellow not first then yellow = 8/11 * 3/10 9/121 + 24/110

OpenStudy (anonymous):

love it :)

OpenStudy (cwrw238):

jake - yes thats the way

OpenStudy (cwrw238):

= 0.29 as a decimal correct to 2 places

OpenStudy (cwrw238):

or as a fraction 177/605